# GeoSDF: Plane Geometry Diagram Synthesis via Signed Distance Field

### v1

@@ 4. Experiments
In this section, we first describe the implementation details, then give results of synthesized diagrams for normal high-school level and IMO level problems in Sec. 4.2 and Sec.4.3, next evaluate the effectiveness of GeoSDF for solving Plane Geometry Problems (PGP) in Sec. 4.4, finally show analysis of our GeoSDF in last three sections.
4.1. Implementation Details
@@
During the optimization process, we set the maximum number of iteration steps to 10,000. The learning rate was annealed from an initial value of 0.1 down to $1\times 10^{-6}$ using a cosine annealing schedule. We define a successful synthesis when the total constraint loss falls below a threshold of 0.1. We select FormalGeo7k (Zhang et al., 2023b) as the source datasets for providing the base geometry predicates as the input constraints of GeoSDF. FormalGeo7k consists of 6,981 PGPs, which were collected and re-annotated from previous PGP datasets, including GeoQA (Chen et al., 2021), GeoQA-Plus (Cao and Xiao, 2022), and Geometry3K (Lu et al., 2021). We leverage the annotated predicates from the FormalGeo7k to synthesize geometry diagrams using GeoSDF by first transforming these predicates into our structural and constraint-based representations and subsequently optimizing the SDF to generate precise and structurally valid geometry diagrams. For the experiments on IMO problems, we use the problems selected by AlphaGeometry (Trinh et al., 2024), and the constraints of these IMO problems were manually annotated by us according to the problem given conditions. In our experiments, we employ an NVIDIA RTX 4090 GPU to accelerate the synthesis of diagrams.
@@
Figure 5. The visualization of the synthesized geometry diagrams. Left: Synthesis on the FormalGeo7k dataset with diversity, in which the first row is the ground truth. Right: Synthesised with detailed annotations.
4.2. Visualization of Synthesized Diagrams
Geometry Diagram Synthesis. We first applied GeoSDF to synthesize geometry diagrams based on the FormalGeo7k dataset. For each problem in the dataset, we leveraged the annotated textual predicates as the problem constraints and then used them to synthesise the corresponding plane geometry diagram. The visualization of five cases of these synthesized diagrams is shown in the left part of Figure 5. For each original problem, we select two new diagrams from the synthesized results, with the top one showing the original diagram (i.e., ground truth) and the other two representing the newly synthesized ones. According to the visualization cases, the synthesized diagrams are visually consistent with the originals but have diverse visual characteristics like rotation, demonstrating GeoSDF’s effectiveness in synthesizing diagrams that have consistent geometric features with original diagrams based on the provided textual predicates. Moreover, the diversity in diagram visual characteristics showcases the GeoSDF’s flexibility and generalization ability, effectively capturing variations while maintaining the underlying geometric relationships. Such flexibility and generalization make GeoSDF a valuable and promising tool for advancing the synthesis and understanding of geometry diagrams in the PGP field.
Geometry Diagram Synthesis with Annotations. Furthermore, our method is highly flexible, providing a range of annotations for synthesized diagrams, including parallelism, perpendicularity, lengths, and angles. These annotations are clearly visualized in Figure 5 (right). This capability could not only enhance human readability but also Multi-Modal LLM understanding of geometric diagrams, guiding models to focus on specific geometric relationships rather than relying solely on abstract visual features, especially since visual encoders may not be optimized for these specific tasks.
Comparison to Existing Methods. A key limitation of prior methods like GeomVerse (Kazemi et al., 2024) and MAVIS (Zhang et al., 2024c) is their reliance on an initial base geometry (e.g., a closed shape like a rectangle or circle) from which other elements are progressively added. In contrast, GeoSDF eliminates this requirement, capable of synthesizing an entire diagram from scratch using only points and lines, without any predefined initial shape.
4.3. International Mathematical Olympiad Geometry
4.3.1. IMO Geometry Diagrams Synthesis and Visualization
A key advantage of GeoSDF is its ability to synthesize complex geometry diagrams, particularly in the context of IMO geometry problems. Previous methods are unable to synthesize these complicated diagrams as they rely on predefined geometric elements and rule-based algorithms to splice these elements together (Kazemi et al., 2024; Deng et al., 2024). These splicing methods are not capable of synthesising IMO problems due to the increased complexity and these methods are not capable to synthesize diagrams based on given conditions, the need for more intricate geometric shapes, more complicated relationships between geometric elements, and a greater number of independent lines that are challenging for previous methods to handle. Therefore, we applied GeoSDF to the FormalGeo-IMO dataset, and in Figure 2 of Section 3 we have shown the visualization of the synthesized geometry diagrams for four different IMO problems. These IMO cases are highly complex, each constructed with different geometry elements and relationships. For example, the diagram in the first problem involves intricate inscribed relationships, the second problem is composed of multiple intersecting line segments, and the third and fourth problems require simultaneously handling the relationships between two circles as well as multiple line segments connected to these circles. The synthesized diagrams by GeoSDF are visually and mathematically consistent with the original IMO problems, demonstrating the robust potential of GeoSDF in handling complex geometry tasks.
4.3.2. Human Evaluation on IMO Geometry Problems
@@
To evaluate whether the synthesized IMO diagrams preserve the geometric characteristics of the original diagram, we conducted a human evaluation study. We invited 20 participants with bachelor’s degrees in science to assess whether the diagrams generated by GeoSDF are structurally equivalent to the original diagram (i.e., ground truth). A total of 30 IMO geometry problems were used in the evaluation. On average, participants judged that 88.67% (26.6 out of 30) of the synthesized diagrams fully reflected the geometry information in the original problem diagram. This result highlights that GeoSDF exhibits strong capability in synthesizing diagrams for IMO-level problems. The design of the questionnaire is included in Appendix LABEL:app:questionnaire.
4.4. Quantifiability of GeoSDF
Figure 6. The example of quantifiability of GeoSDF. Our method is able to measure all the values in the diagram (e.g. degree, area), where the left diagram is the original diagram and the right one is synthesised by GeoSDF.
@@ 4.4. Quantifiability of GeoSDF
One of the significant property of GeoSDF is the quantifiability. It means that all elements of the geometry diagram can be directly quantified or measured (e.g. degree of angle) and extracted (area and length will be scaled through the scale bar of each case). For example, the first problem in Fig. 6 asks for the measure of angle CDE (marked by the red arrow). Once the SDF optimization is complete and the diagram meets the loss constraint, we extract the angle value as the solution, which is 55. Similarly, in another example from Fig. 6, we determine the length of line YZ, the diameter of circle M. This capability enables us to extend its application to a broader range of experiments, such as direct problem solving and verifying whether the synthesized diagrams precisely satisfy the original problem.
4.4.1. Plane Geometry Problem Solving
In this way, we conduct the plane geometry problem (PGP) solving task on the GeoQA (Chen et al., 2021) dataset and comparing with various SOTA solvers includes general neural-based and MLLM-based methods. Specifically, we first synthesize the SDF representation of the problem by constructing it from the symbolic clauses given by the problem. We then measure the values of the target elements in the problem, treating them as the solution. This approach corresponds to the completion setting, where the solution value is directly output. The experiment results are shown in Table 2. Surprisingly, due to its high accuracy in diagram synthesis, our method achieved a 94.5% accuracy rate under the completion setting, significantly outperforming all current SOTA solvers, including both neural-based and MLLM-based PGP solvers. These results demonstrate the strong potential of our approach for solving PGPs. In the Choice setting, our method also significantly outperforms other models, including multi-modal LLMs equipped with Chain-of-Thought reasoning or specifically fine-tuned for PGP solving tasks. Notably, our method relies solely on the basic problem information parsed from the original input using FormalGeo, without incorporating any additional data which ensuring a fair comparison with other models. These experiments highlight the strong potential of applying GeoSDF in the field of PGP solving research. GeoSDF also shows promise as a practical tool to support the development of PGP solvers, like serving as an external module for solution verification or diagram-based reasoning.
@@ 4.4.1. Plane Geometry Problem Solving
Table 2. Accuracy (%) on the GeoQA test set and reported for two settings: Completion, the answer value is directly outputted; Choice, the problem choice is directly outputted. For neural PGP solvers, a random choice is selected if the completion result doesn’t match any given choice.
@@
NGS(Chen et al., 2021) 60.0 69.5
DPE-NGS(Cao and Xiao, 2022) 62.7 70.4
SCA-GPS(Ning et al., 2023) 64.1 72.7
@@
G-LLaVA-13B(Gao et al., 2025) - 67.0
GeoX(Xia et al., 2025) 54.9 -
Qwen2.5-VL-7B(Bai et al., 2025) 64.2 64.3
InternVL2.5-8B(Chen et al., 2024) 50.5 58.1
MAVIS-7B(Zhang et al., 2024c) - 68.3
R-CoT-8B(Deng et al., 2024) - 75.1
GeoSDF-Solving 94.5 95.9
4.4.2. Synthesised Diagram Structural Equivalence Check
@@
In addition, when using GeoSDF to synthesize diagrams based on known problem statements, questions, and answers, the quantifiability of GeoSDF enables further verification of whether the synthesized diagrams satisfy the given conditions and correspond to the correct answers. If the values extracted from the SDF representation align with the expected solution and the optimization loss remains below a predefined threshold, we consider the diagram a faithful representation of the original problem. This verification process uses the problem’s goal formula to locate and extract relevant geometric features from the SDF object. To quantify this, we evaluated the success rate of diagram synthesis on the FormalGeo7K dataset. Among 6,981 samples, 5,943 diagrams (85.13%) passed the loss constraint check, and 5,697 diagrams (81.60%) successfully meet the target equivalence conditions.
4.5. Equivalence of Synthesised GeoSDF Diagrams
To further assess whether the diagrams synthesized by GeoSDF preserve the same geometry information in the original diagrams, we conducted an experiment by replacing the geometry diagrams in the test set of GeoQA with newly synthesized versions. We then evaluated whether baseline models could still solve the problems using these new diagrams. We tested several baseline models, including NGS (Chen et al., 2021) and G-LLaVA-7B (Gao et al., 2025), using both the original and GeoSDF-synthesized test sets. All models were evaluated with their publicly released code and model weights were trained on the original training set. To create the new diagrams, we annotated problem predicates and used GeoSDF to synthesize diagrams for each test example. The results of this experiment are shown in Table 3. Interestingly, the accuracy on the GeoSDF-synthesized test set was slightly higher than on the original test set. We attribute this improvement to the fact that some of the original diagrams were of relatively low quality, which likely hindered model performance. In contrast, the diagrams synthesized by GeoSDF offer clearer visual representations, helping the models better understand the geometry and ultimately improving accuracy.
@@ 4.5. Equivalence of Synthesised GeoSDF Diagrams
Table 3. Comparison (%) on the GeoQA test set between original diagrams and those synthesized by GeoSDF, evaluated across different base models.
Test Set Original Test Set Synthesised by GeoSDF
NGS (Chen et al., 2021) 60.9 61.5
SCA-GPS(Ning et al., 2023) 64.1 64.9
G-LLaVA-7B(Gao et al., 2025) 64.2 64.3
4.6. Computational Efficiency and Convergence Analysis
The computational efficiency and convergence behavior of our GeoSDF framework are critical for its practical applicability. To evaluate computational performance, we compare the performance, resource usage, and rendering time versus Batch Size for the IMO level figure as shown in Tab. 4. GeoSDF could also run on the CPU or the laptop platform due to it is not sensitive to the computing power. The optimization time is nearly constant, so not included in this table because we find the bottleneck is in the transfer speed between the GPU and CPU. We measure the average optimization time for generating diverse geometric figures. For typical complex IMO level figures (e.g., those involving 20+ elements and 20+ constraints), the generation process converges in about 20s with annotations. For example, the optimization time is about 21.3s for IMO 2021 P4 for all batch size configurations with 24 elements and 24 constraints. This performance demonstrates the feasibility of our approach for interactive or near-interactive geometric design.
@@ 4.6. Computational Efficiency and Convergence Analysis
Table 4. Performance, Resource Usage, and Rendering Time vs. Batch Size over IMO Diagram Synthesis. The accuracy is calculated as the loss less than 0.1 in a batch. The Memory is the maximum GPU memory allocated throughout the entire process.
Batch Size Accuracy (%) Memory (GB) Rendering Time (s)
2048 57.67 14.49 77.4
1024 57.42 7.02 39.2
512 56.84 3.52 21.4
256 60.16 1.77 12.3
128 60.94 0.89 7.5
64 62.19 0.50 5.2
32 53.13 0.38 4.0
4.7. Impact of Crowd Regularization

### v2

@@ 4 Experiments
In this section, we first describe the implementation details, then give results of synthesized diagrams for normal high-school level and IMO level problems in Sec. 4.2, next evaluate the effectiveness of GeoSDF for solving Plane Geometry Problems (PGP) in Sec. 4.4, finally show analysis of our GeoSDF in the last three sections.
4.1 Implementation Details
Our framework is implemented in Python using the PyTorch library and is structured as a user-friendly package for public release. For the optimization, we use the AdamW optimizer for a maximum of 10,000 iterations. The learning rate follows a cosine annealing schedule, decaying from an initial value of $0.1$ down to $1\times 10^{-6}$. A synthesis is deemed successful if the total constraint loss falls below a threshold of 0.1.
@@ 4.1 Implementation Details
In our experiments, we select FormalGeo7k Zhang et al. (2023b) as the source datasets for providing the base geometry predicates as the input constraints of GeoSDF. FormalGeo7k consists of 6,981 PGPs, which were collected and re-annotated from previous PGP datasets, including GeoQA Chen et al. (2021), GeoQA-Plus Cao & Xiao (2022), and Geometry3K Lu et al. (2021). We leverage the annotated predicates from the FormalGeo7k to synthesize geometry diagrams using GeoSDF by first transforming these predicates into our structural and constraint-based representations and subsequently optimizing the SDF to generate precise and structurally valid geometry diagrams. For more challenging problems, we use the IMO problems selected by AlphaGeometry Trinh et al. (2024), for which we manually annotated the constraints from the problem statements.
@@
Figure 3: Qualitative results. Left: Synthesis on the FormalGeo7k dataset with diversity yet geometrically consistent (first row: ground truth). Middle: Synthesised with detailed annotations. Right: Synthesised IMO geometric problems with annotations.
4.2 Qualitative Synthesis Results
Geometry Diagram Synthesis. We first applied GeoSDF to synthesize geometry diagrams based on the FormalGeo7k dataset. For each problem in the dataset, we leveraged the annotated textual predicates as the problem constraints and then used them to synthesise the corresponding plane geometry diagram. To demonstrate our method’s ability to generate diverse outputs, we augmented the FormalGeo7k dataset by synthesizing 32 diagrams with diversity for each problem description from 7k to 224k diagrams. This process significantly increases the visual variety available for training and evaluation. A random selection of 400 synthesized examples is provided in the Appendix.
The visualization of five cases of these synthesized diagrams is shown in the left part of Figure 3. For each original problem, we select two new diagrams from the synthesized results, with the first row showing the ground truth and the other two representing the newly synthesized ones. The synthesized diagrams are geometrically consistent with the originals, but with rotation, mirror, and deformation diversity. Furthermore, our method provides a range of annotations for synthesized diagrams to enhance human readability, including parallelism, perpendicularity, lengths, areas, and angles. These annotations are clearly visualized in Figure 3 (Middle and Right).
IMO Geometry Diagrams Synthesis. Our GeoSDF can synthesize geometrically complex diagrams, particularly those found in International Mathematical Olympiad (IMO) problems. Previous methods, limited by their reliance on rule-based algorithms, which unable to generate such diagrams. Meanwhile, the model-based methods struggle with the deterministic rules and mathematical validity, resulting in synthesizing inaccurate diagrams. To demonstrate our method’s ability to handle these challenges, we visualize the synthesis of IMO problems (from 1959 to 2022) in Figure 3 (right). These examples showcase our framework’s ability to deal with diverse and challenging constraints. The resulting diagrams are both visually coherent and mathematically precise, demonstrating the potential of GeoSDF in handling sophisticated geometry synthesis tasks.
@@ 4.2 Qualitative Synthesis Results
To verify the correctness of our synthesized IMO diagrams, we conducted a human evaluation study. We invited 20 participants with bachelor’s degrees in science to assess whether the diagrams generated by GeoSDF are geometrically equivalent to the original diagram (i.e., ground truth). A total of 30 IMO geometry problems were used in the evaluation. On average, participants judged that 88.67% (26.6 out of 30) of the synthesized diagrams fully reflected the geometry information in the original problem diagram. This result highlights that GeoSDF exhibits strong capability in synthesizing diagrams for IMO-level problems. The design of the questionnaire is included in Appendix G.
4.3 Directly Generating Diagrams from Natural Language
We assess the natural-language–to-diagram stage by evaluating the parser that converts natural problem text into symbolic constraints. Constraints are compared as order-independent sets, with F1 and Jaccard similarity used to measure overlap while balancing false positives and negatives Manning (2008). Before scoring, we apply light preprocessing: standardizing text (e.g., lowercasing, sorting arguments) and tokenizing while preserving expressions such as Collinear(A,B,C). All evaluations are conducted on the GeoQAChen et al. (2021) test set. The parser achieves solid overall performance with an F1 of 87.74% and Jaccard of 83.53%, showing that most constraints are captured reliably. These results suggest that the symbolic representation produced from natural text is able to support subsequent SDF-based optimization. While occasional parsing errors remain, they rarely alter the overall structure of the diagram, indicating the feasibility of the pipeline. We further provide a complete parsing example in Appendix F.
4.4 Quantifiability of GeoSDF
@@
One of the significant properties of GeoSDF is the quantifiability. It means that all elements of the geometry diagram can be directly quantified or measured (e.g. degree of angle) and extracted (area and length will be scaled through the scale bar of each case). For example, the first problem in Fig. 4 asks for the measure of angle CDE (marked by the red arrow). Once the SDF optimization is complete and the diagram meets the loss constraint, we extract the angle value as the solution, which is 55. Similarly, in another example from Fig. 4, we determine the length of line YZ, the diameter of circle M. This capability enables us to extend its application to a broader range of experiments, such as direct problem solving and verifying whether the synthesized diagrams precisely satisfy the original problem.
Figure 4: The example of quantifiability of GeoSDF. Our method is able to measure all the values in the diagram (e.g. degree, area), where the left diagram is the original diagram and the right one is synthesised by GeoSDF.
4.4.1 Plane Geometry Problem Solving
A key application of our framework is solving Plane Geometry Problems (PGPs) through a novel ”solve-by-construction” paradigm. GeoSDF finds solutions by first synthesizing a high-fidelity diagram that satisfies all given geometric constraints and then directly measuring the properties of the target element (e.g., an angle’s degree or a segment’s length) from the final SDF representation.
We evaluated on the GeoQA Chen et al. (2021) dataset, comparing its performance against SOTA neural and MLLM-based solvers. To provide a comprehensive analysis, we test our method in three distinct settings, with results presented in Table 2. GeoSDF-GT (Oracle Input): This is our main result. We use the ground-truth symbolic constraints provided in the dataset to synthesize the diagram. This setting isolates the performance of the core synthesis and measurement pipeline, removing any noise from upstream parsing errors. GeoSDF (End-to-End): To simulate a real-world scenario, this setup uses a fine-tuned LLM to parse the problem’s text and diagram image into symbolic constraints, which are then fed into our solver. This evaluates the entire end-to-end pipeline.
Our primary method, GeoSDF, achieves a remarkable 94.5% accuracy in the completion setting and 95.9% in the choice setting. This result significantly surpasses all existing solvers, which perform logical or numerical reasoning. This highlights the potential of the solve-by-construction approach for a large part of geometry problems where the answer is a measurable quantity. Even in the end-to-end setting (GeoSDF), our method achieves SOTA accuracy of 78.5%, demonstrating its practical viability despite the challenges of automated parsing. These experiments underscore the power of high-precision diagram synthesis as a direct path to problem-solving. Furthermore, GeoSDF’s quantifiable nature makes it a promising tool for external solution verification or as a reasoning module within hybrid PGP solvers. A detailed analysis of failure cases is provided in Appendix D.
@@ 4.4.1 Plane Geometry Problem Solving
Table 2: Accuracy (%) on the GeoQA test set, the results are reported for two settings: Completion, the answer value is directly outputted; Choice, the problem choice is directly outputted. For neural PGP solvers, a random choice is selected if the completion result doesn’t match any given choice.
@@
NGSChen et al. (2021) 60.0 69.5
DPE-NGSCao & Xiao (2022) 62.7 70.4
SCA-GPSNing et al. (2023) 64.1 72.7
@@
G-LLaVA-13BGao et al. (2025) - 67.0
GeoXXia et al. (2025) 54.9 -
Qwen2.5-VL-7BBai et al. (2025) 64.2 64.3
InternVL2.5-8BChen et al. (2024) 50.5 58.1
MAVIS-7BZhang et al. (2025) - 68.3
R-CoT-8BDeng et al. (2024) - 75.1
GeoGen-SFT-7BPan et al. (2025) 64.6 78.0
GeoUni-1.5BCheng et al. (2025) 66.7 78.0
GeoSDF 78.5 83.2
GeoSDF-GT 94.5 95.9
4.4.2 Synthesised Diagram Structural Equivalence Check
@@
When using GeoSDF to synthesize diagrams based on known problem statements, questions, and answers, the quantifiability of GeoSDF enables further verification of whether the synthesized diagrams satisfy the given conditions and correspond to the correct answers. If the values extracted from the SDF representation align with the expected solution and the optimization loss remains below a predefined threshold, we consider the diagram a faithful representation of the original problem. This verification process uses the problem’s goal formula to locate and extract relevant geometric features from the SDF object. To quantify this, we evaluated the success rate of diagram synthesis on the FormalGeo7K dataset. Among 6,981 samples, 5,943 diagrams (85.13%) passed the loss constraint check, and 5,697 diagrams (81.60%) successfully meet the target equivalence conditions. Upon evaluating the dataset, we identified that the main source of synthesis failures involves problems that are not strictly value-based. These cases typically include undefined variables within their geometric constraints, such as defining a line segment’s length as ’x+2’. While our parsing module capably translates these statements into symbolic constraints, the GeoSDF synthesis process cannot render a diagram without concrete numerical values, leading to a failure in visualization.
4.5 Equivalence of Synthesised GeoSDF Diagrams
@@
Table 3: Comparison (%) on the GeoQA test set between original diagrams and those synthesized by GeoSDF, evaluated across different base models.
Test Set
Original
Test Set
Synthesised
by GeoSDF
NGS Chen et al. (2021) 60.9 61.5
SCA-GPSNing et al. (2023) 64.1 64.9
G-LLaVA-7BGao et al. (2025) 64.2 64.3
To further assess whether the diagrams synthesized by GeoSDF preserve the same geometry information in the original diagrams, we conducted an experiment by replacing the geometry diagrams in the test set of GeoQA with newly synthesized versions. We then evaluated whether baseline models could still solve the problems using these new diagrams. We tested several baseline models, including NGS Chen et al. (2021) and G-LLaVA-7B Gao et al. (2025), using both the original and GeoSDF-synthesized test sets. All models were evaluated with their publicly released code, and model weights were trained on the original training set. To create the new diagrams, we annotated problem predicates and used GeoSDF to synthesize diagrams for each test example. The results of this experiment are shown in Table 3. Interestingly, the accuracy on the GeoSDF-synthesized test set was slightly higher than on the original test set. We attribute this improvement to the fact that some of the original diagrams were of relatively low quality, which likely hindered model performance. In contrast, the diagrams synthesized by GeoSDF offer clearer visual representations, helping the models better understand the geometry and ultimately improving accuracy.
4.6 Computational Efficiency and Convergence Analysis
@@
Table 4: Performance, Resource Usage, and Time vs. Batch Size over IMO Diagram Synthesis. The Successful Rate is calculated as the loss less than 0.1 in a batch. The Memory is the maximum GPU memory allocated throughout the entire process. We take 10000 steps for optimization.
Batch Size Successful Rate (%) Memory (GB) Time (s)
2048 57.67 14.49 98.7
1024 57.42 7.02 60.5
512 56.84 3.52 42.7
256 60.16 1.77 33.6
128 60.94 0.89 28.8
64 62.19 0.50 26.5
32 53.13 0.38 25.3
The computational efficiency and convergence behavior of our GeoSDF framework are critical for its practical applicability. To evaluate computational performance, we compare the performance, resource usage, and time versus Batch Size for the IMO level figure as shown in Tab. 4. GeoSDF could also run on the CPU or the laptop platform due to it is not sensitive to the computing power. The optimization time is nearly constant versus Batch Size, and we find the bottleneck is in the transfer speed between the GPU and CPU. We measure the average optimization time for generating diverse geometric figures. For typical complex IMO level figures (e.g., those involving 20+ elements and 20+ constraints), the generation process converges in about 20s with annotations. For example, the optimization time is about 21.3s for IMO 2021 P4 for all batch size configurations with 24 elements and 24 constraints. This performance demonstrates the feasibility of our approach for interactive or near-interactive geometric design.
4.7 Impact of Crowd Regularization